January 2009 Julius purchased for a hundred and ten thousand a rental house which he made Italy placed in service. The purchase price included $10,000 for the land in August of 2018, he decided to sell the house to the tenant how much depreciation may he deduct in 2018
Answer #1

According to the rules laid down by the IRS,

To find the depreciation, there can be two ways:

  1. Take a cost – basis and divide it by 27.5 (agreed upon estimated life by IRS).
  2. Depreciate 3.636% of the cost basis of the rental property each year.

SOLUTION:

Annual depreciation to be charged for 2009 to 2017

= (Cost basis * 3.636%) * 9 years

= (100,000 * 3.636%) * 9

= $3,636 * 9

= $32,724

*(Cost basis doesn’t include the value of land.)

Depreciation for the year 2018 (up to August)

= 100,000 * 2.273% *(reason is that it was used for only 7 months and its solved acc. To IRS table)

= $2273

Hence, TOTAL DEPRECIATION CHARGEABLE:

= 32,724 + 2273

= $ 34,997